### v1

# Vector Quantized-Elites: Unsupervised and Problem-Agnostic Quality-Diversity Optimization

Abstract
Quality-Diversity algorithms have transformed optimization by prioritizing the discovery of diverse, high-performing solutions over a single optimal result. However, traditional Quality-Diversity methods, such as MAP-Elites, rely heavily on predefined behavioral descriptors and complete prior knowledge of the task to define the behavioral space grid, limiting their flexibility and applicability. In this work, we introduce Vector Quantized-Elites (VQ-Elites), a novel Quality-Diversity algorithm that autonomously constructs a structured behavioral space grid using unsupervised learning, eliminating the need for prior task-specific knowledge. At the core of VQ-Elites is the integration of Vector Quantized Variational Autoencoders, which enables the dynamic learning of behavioral descriptors and the generation of a structured, rather than unstructured, behavioral space grid — a significant advancement over existing unsupervised Quality-Diversity approaches. This design establishes VQ-Elites as a flexible, robust, and task-agnostic optimization framework. To further enhance the performance of unsupervised Quality-Diversity algorithms, we introduce two key components: behavioral space bounding and cooperation mechanisms, which significantly improve convergence and performance. We validate VQ-Elites on robotic arm pose-reaching and mobile robot space-covering tasks. The results demonstrate its ability to efficiently generate diverse, high-quality solutions, emphasizing its adaptability, scalability, robustness to hyperparameters, and potential to extend Quality-Diversity optimization to complex, previously inaccessible domains.
Index Terms:
Quality-Diversity, Unsupervised Learning, MAP-Elites, Autonomous Robots
I Introduction
Figure 1: Overview of VQ-Elites
Traditional evolutionary algorithms (EAs) have been a cornerstone of optimization, designed primarily to identify a single optimal solution for a given problem. Although effective for well-defined tasks, this narrow focus on singular objectives limits their ability to address complex challenges, such as those in robotics and autonomous systems. This limitation hinders their adaptability in dynamically changing environments, where a diverse range of high-performing solutions is often required. Consequently, traditional EAs struggle in scenarios that demand versatility rather than a single, best solution [1, 2].
Quality-Diversity (QD) algorithms address the limitations of traditional evolutionary algorithms by redefining the optimization objective [1, 2]. Rather than seeking a single optimal solution, QD algorithms, such as MAP-Elites [3, 4], aim to uncover a diverse set of high-performing solutions across a broad behavioral landscape. This “illumination” approach generates multiple effective parameter vectors, each associated with a unique behavior. By promoting both diversity and quality, QD algorithms enable more flexible and robust applications, particularly in autonomous systems, where a repertoire of high-quality behaviors is essential for adapting to dynamic challenges. For instance, QD algorithms have demonstrated the ability to generate diverse locomotion behaviors in quadruped robots [4, 5, 6].
Despite these advances, QD algorithms face significant limitations, particularly regarding the reliance on predefined or ”hard-coded” behavioral descriptors (BDs). In MAP-Elites, for example, the behavior space is often discretized using a set of user-defined descriptors to categorize behaviors, which imposes constraints on the algorithm’s adaptability. These hard-coded descriptors assume a priori knowledge about the task or environment and may not fully capture the complexity or diversity inherent in high-dimensional, novel tasks. This constraint limits the algorithm’s flexibility, as predefined descriptors may inadvertently filter out behaviors that fall outside of the original descriptors or misrepresent complex solutions. Moreover, for tasks where the behavior descriptors are hard to design or even unknown, the manual design of BDs becomes impractical and can hinder the QD algorithm’s capacity to scale or generalize to new domains.
Early approaches to automatic BD extraction relied on large static datasets to learn lower-dimensional representations using techniques like Principal Component Analysis (PCA) or auto-encoders. For instance, Cully and Demiris [7] utilized a dataset of handwritten digits to train an auto-encoder, creating a low-dimensional latent space that captured the key features of different digits. This latent space then served as the behavioral descriptor space, enabling QD algorithms to generate a diverse set of solutions spanning the learned feature space. This approach allowed a robotic arm to autonomously learn to draw digits without requiring manual definitions of their key characteristics.
The challenges posed by predefined BDs and static datasets encouraged the development of unsupervised learning methods for BD extraction. A notable example is the AURORA algorithm [8, 9], which employs neural network-based representation learning to autonomously identify behavior descriptors directly from the solution space. The main innovation compared to using a static dataset is that AURORA uses the data generated by the QD algorithm to train the auto-encoder in an online fashion. This innovation represented a significant leap, allowing QD algorithms to generalize across a broader range of tasks without relying on hand-crafted descriptors, thus enabling more adaptable and versatile QD frameworks. Nevertheless, AURORA still requires a lot of heuristics and extensive fine-tuning to be performant.
In this paper, we introduce Vector Quantized-Elites (VQ-Elites), a novel QD algorithm that uses unsupervised learning to autonomously construct meaningful behavior spaces without prior task or environment knowledge (Fig. 1). We leverage a Vector Quantized Variational Autoencoder (VQ-VAE) to make our algorithm able to dynamically learn behavior descriptors, thus enabling autonomous discovery of relevant skills and behaviors. By integrating VQ-VAE into the QD framework, our approach organizes the behavior space into a structured grid by automatically clustering latent representations. This grid-based structure addresses the shortcomings of previous unsupervised BD methods by combining flexibility with the advantages of well-structured archives. Moreover, VQ-Elites enhances the scalability of QD algorithms, facilitating exploration in high-dimensional tasks and environments that traditional QD methods struggle to handle.
To demonstrate the effectiveness of VQ-Elites, we evaluate its performance across a variety of tasks that highlight its adaptability and robustness in complex scenarios. These include both structured and unstructured space-covering tasks, as well as high-dimensional pose-reaching tasks for robotic manipulators. In each experiment, we assess VQ-Elites’ ability to generate diverse, high-quality behaviors while maintaining an organized archive. The results underscore its strengths in flexibility and robustness, showcasing its potential to advance QD frameworks for autonomous robotic systems and extend their applicability to more complex problem domains.
In short:
  • We propose VQ-Elites, a novel QD algorithm that integrates unsupervised learning to construct a flexible, task-agnostic behavior space. Using VQ-VAE models, VQ-Elites dynamically learns BDs, enabling the autonomous discovery of relevant behaviors across diverse tasks without prior environment-specific knowledge.
  • VQ-Elites autonomously creates a structured grid in behavior space by clustering latent representations, addressing the limitations of previous unsupervised BD methods while maintaining the benefits of structured archives.
  • We introduce two components, Behavioral Space Bounding and Cooperation, which drastically improve the convergence of unsupervised QD algorithms.
  • We evaluate VQ-Elites across diverse tasks, including space-covering challenges, pose-reaching tasks for robotic manipulators, and scenarios involving high-dimensional data, and constrained systems.
  • Experimental results show that VQ-Elites efficiently generates diverse, high-quality behaviors while preserving a well-organized archive, highlighting its potential to enhance QD frameworks for autonomous robotic systems and extend their applicability to complex problem domains.
II Preliminaries
II-A Quality-Diversity Algorithms
One of the notable strengths of QD algorithms is their ability to provide multiple high-performing solutions, offering users significant flexibility. This flexibility arises from the wide range of solutions available, allowing users to balance complexity, desired behavior, and practical implementation.
To express this more formally, consider an objective function that returns a fitness value, $f_{\theta}$, and a behavioral descriptor vector, $\boldsymbol{b}_{\theta}$, which defines the behavior:
$f_{\theta},\boldsymbol{b}_{\theta}\leftarrow f(\boldsymbol{\theta})$ (1)
In practice, the behavioral descriptor characterizes the solution, while the fitness value measures its quality. To clarify the concept of a BD and the difference to the fitness, consider a scenario where a robot must traverse various locations within an environment. In this case, the robot’s coordinates can be the BD, while the average speed it reaches the targets the fitness value. Bringing these components together, the optimization algorithm generates multiple solutions, each defined by its BD vector, which describes the behavior or outcome, and its fitness value, which quantifies its effectiveness.
But, what is our ultimate goal when using a Quality-Diversity algorithm? Let us assume, without the loss of generality, that we have a behavior space (feature space) $\mathcal{B}$ and a fitness function that we want to maximize. The ultimate goal of QD optimization is to find all parameters, $\boldsymbol{\theta}$, that maximize the fitness function for every $\boldsymbol{b}\in\mathcal{B}$:
$\forall\boldsymbol{b}\in\mathcal{B}\qquad\boldsymbol{\theta}^{*}=\operatorname*{arg\,max}_{\boldsymbol{\theta}}f_{\theta}$
$s.t.\qquad\boldsymbol{b}=\boldsymbol{b_{\theta}}$ (2)
As already mentioned, the result of a QD algorithm is a set of solutions. This set is called either the archive, collection, or map. Furthermore, during the optimization process, it is possible that the algorithm can produce a solution with a similar BD to a solution that already exists in the archive. In that case, and since we cannot keep all possible solutions, the two solutions compete, and finally, the fittest one is preserved.
Archives in QD algorithms can take various forms. Perphaps the easiest one involves discretizing the behavior space into a grid, where each cell corresponds to a unique solution. Solutions generated by the algorithm are assigned to the closest cell based on the Euclidean distance between the solution’s BD and the cell’s center. This grid-based method is widely used, particularly in the MAP-Elites algorithm, one of the most popular QD optimization approaches. Other archive variations include more abstract and unstructured designs. For instance, a user may define a distance threshold, allowing a solution to be stored if its distance to all existing BDs exceeds the threshold, or if it outperforms competing BDs within the threshold.
Algorithm 1 MAP-Elites algorithm
procedure MAP-Elites($[n_{1},...,n_{d}]$)
     $\mathcal{A}\longleftarrow$ create_empty_archive($[n_{1},...,n_{d}]$)
     for $i=1\to G$ do $\triangleright$ Initialization: $G$ random $\boldsymbol{\theta}$
         $\boldsymbol{\theta}=$ random_solution()
         add_to_archive($\boldsymbol{\theta},\mathcal{A}$)
     end for
     for $i=1\to I$ do $\triangleright$ Main loop, $I$ iterations
         $\boldsymbol{\theta}=$ selection($\mathcal{A}$)
         $\boldsymbol{\theta}^{\prime}=$ variation($\boldsymbol{\theta}$)
         add_to_archive($\boldsymbol{\theta}^{\prime},\mathcal{A}$)
     end for
     return $\mathcal{A}$
end procedure
procedure add_to_archive($\boldsymbol{\theta},\mathcal{A}$)
     $(p,\mathbf{b})\longleftarrow$ evaluate($\boldsymbol{\theta}$)
     $c\longleftarrow$ get_cell_index($\mathbf{b}$)
     if $\mathcal{A}(c)=null$ or $\mathcal{A}(c).p<p$ then
         $\mathcal{A}(c)\longleftarrow p,\boldsymbol{\theta}$
     end if
end procedure
As a reference, we provide the pseudo-code of the MAP-Elites algorithm (Algo. 1), one of the first and most used QD algorithms. The popularity of MAP-Elites as a QD algorithm stems from its simplicity, both in terms of understanding and implementation. However, its main challenge lies in discretizing the behavior space, which is highly problem- and domain-dependent. This becomes particularly difficult in scenarios involving high-dimensional features or when the behavior space lacks well-defined constraints.
II-B Unsupervised Representation Learning
Representation learning, is a fundamental concept in machine learning and artificial intelligence, focusing on the development of effective data representations that enhance the performance of machine learning models on various tasks. The goal is to automatically transform raw data, such as images, text, or audio, into a more abstract and useful format that captures the essential features of the data while discarding irrelevant or redundant information. Effective representations can simplify learning tasks, reduce the need for domain-specific feature engineering, and enable generalization across different datasets and tasks. In contrast, traditional approaches relied heavily on manual feature engineering, requiring domain experts to identify and extract useful features. This process, however, is often labor-intensive, prone to biases, and limits scalability across different problem domains. Representation learning automates this process, allowing models to discover and learn optimal features from data autonomously. This paradigm shift has led to breakthroughs in various fields, including natural language processing [10], computer vision [11, 12], and speech recognition [13, 14].
VQ-VAE represents a powerful extension of traditional autoencoders, combining ideas from variational autoencoders (VAEs) and vector quantization. Introduced by Oord et al. [15], VQ-VAE addresses the limitations of continuous latent spaces by discretizing them, allowing the model to learn discrete latent representations that are often more interpretable and suitable for tasks such as generation, compression, and unsupervised learning. VQ-VAE has shown remarkable performance in applications such as image and audio synthesis, video generation, and reinforcement learning [16, 17].
VQ-VAE introduces a discrete latent space by using vector quantization. The encoder maps the input $\boldsymbol{x}$ to a latent representation, but instead of sampling from a continuous distribution, the latent code is quantized by selecting the closest code from a predefined set of embeddings, known as the codebook. Let $\mathcal{C}\in\mathbb{R}^{K\times D}$ represent the codebook, where $K$ is the number of embeddings and $D$ is the dimensionality of each embedding vector. For each input $\boldsymbol{x}$, the encoder produces a latent vector $\boldsymbol{z}_{e}(\boldsymbol{x})$, and the quantized vector $\boldsymbol{z}_{q}(\boldsymbol{x})$ is obtained by choosing the nearest codebook entry using the following operation:
$\boldsymbol{z}_{q}(\boldsymbol{x})=\arg\min_{\boldsymbol{c}_{i}\in\mathcal{C}}\|\boldsymbol{z}_{e}(\boldsymbol{x})-\boldsymbol{c}_{i}\|^{2}$ (3)
This step discretizes the latent space, as the latent vector $\boldsymbol{z}_{q}(\boldsymbol{x})$ is always chosen from a finite set of possible embeddings $\{\boldsymbol{c}_{1},\boldsymbol{c}_{2},\dots,\boldsymbol{c}_{K}\}$. The training of a VQ-VAE is performed through back-propagation with the goal of minimizing a loss function. The overall loss function of VQ-VAE is composed of three key components:
  • Reconstruction Loss: This term ensures that the decoder reconstructs the input data accurately from the quantized latent code $\boldsymbol{z}_{q}(\boldsymbol{x})$. Similar to traditional autoencoders, the reconstruction loss is typically measured as the $L_{2}$-norm between the input and the reconstructed output $\hat{\boldsymbol{x}}$:
    $\mathcal{L}_{\text{recon}}=\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|^{2}$ (4)
  • Codebook Loss (Commitment Loss): Since the latent space is quantized, the encoder is encouraged to produce latent vectors that are close to the nearest codebook vectors. This is done by adding a commitment loss that penalizes large differences between $\boldsymbol{z}_{e}(\boldsymbol{x})$ and $\boldsymbol{z}_{q}(\boldsymbol{x})$. The commitment loss pushes the codebook vectors to move towards the encoded latent vectors, ensuring that the quantized vectors are close representations of the original latent vectors:
    $\mathcal{L}_{\text{codebook}}=\|\texttt{sg}[\boldsymbol{z}_{e}(\boldsymbol{x})]-\boldsymbol{c}\|^{2}$ (5)
    where $\texttt{sg}[\cdot]$ represents the stop-gradient operator, which prevents gradients from being back-propagated through the codebook vectors.
  • Commitment Loss for Encoder: To prevent the encoder from fluctuating too far from the chosen codebook vectors, a commitment loss is also applied to the encoder:
    $\mathcal{L}_{\text{commit}}=\beta\|\boldsymbol{z}_{e}(\boldsymbol{x})-\texttt{sg}[\boldsymbol{c}]\|^{2}$ (6)
    Here, $\beta$ is a hyperparameter that controls the trade-off between minimizing reconstruction error and ensuring the encoder’s output aligns with the quantized latent space.
Thus, the total VQ-VAE loss is given by:
$\mathcal{L}_{\text{VQ-VAE}}=\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{codebook}}+\mathcal{L}_{\text{commit}}$ (7)
One of the key challenges of minimizing this combined loss is the non-differentiability of the quantization step (i.e., selecting the nearest codebook entry), which prevents gradients from flowing back through the encoder during standard training. To address this, VQ-VAE uses a straight-through estimator for gradient updates. Specifically, during the backward pass, the gradients from the decoder are passed through the quantization operation unchanged, which allows effective training despite the discrete nature of the latent space [15].
III Related Work
The MAP-Elites algorithm, introduced by Mouret and Clune [3], established a new paradigm in evolutionary optimization by illuminating diverse regions of the solution space rather than focusing solely on a single optimal solution. MAP-Elites partitions the behavioral space into user-defined niches and searches for the highest-performing solution within each niche. This approach has shown success in robotic control tasks, constrained optimization, and applications requiring broad behavioral diversity [4]. However, MAP-Elites relies heavily on pre-defined behavioral descriptors (BDs), which limits its flexibility and adaptability to dynamic or complex environments where optimal descriptors are not readily apparent.
To address this limitation, various adaptations have been proposed. The CVT-MAP-Elites algorithm uses centroidal Voronoi tessellations (CVTs) to partition the behavior space based on well-spread centroids, allowing for more natural, high-dimensional behavior mappings without rigid grid structures [18]. Despite its effectiveness in certain applications, CVT-MAP-Elites still depends on predefined centroids, which may not capture the full range of possible behaviors in a dynamic environment.
More recent work has explored integrating reinforcement learning (RL) techniques with MAP-Elites to improve search efficiency and expand behavioral diversity. For example, methods like PGA-MAP-Elites and QD-PG leverage policy gradients alongside MAP-Elites’ mutation-based search, enabling gradient-informed exploration of high-dimensional behaviors [19, 20]. However, these approaches still rely on predefined descriptors or additional complexity from combining RL methods, which may not be ideal for fully unsupervised BD discovery.
In parallel, advancements in skill discovery through hierarchical reinforcement learning (HRL) and skill chaining [21, 22] have shown the importance of decomposing tasks into reusable sub-skills. These methods provide a framework for discovering hierarchical behaviors but often require substantial supervision or domain knowledge. Unsupervised skill discovery methods, such as DIAYN (Diversity is All You Need) [23] and VALOR [24], emphasize maximizing diversity in the skill space while learning useful policies. However, these approaches primarily focus on policy-level diversity without leveraging structured archives like MAP-Elites for comprehensive exploration.
The introduction of unsupervised learning methods into the QD framework, like AURORA [9], has enabled the automatic extracting latent behavior representations, allowing QD algorithms to operate independently of pre-defined descriptors. For instance, AURORA employs neural networks to identify meaningful patterns within the behavior space, yet it lacks a structured archive organization akin to MAP-Elites, which can hinder systematic exploration in complex domains [9] (as we will show in our experiments). The TAXONS algorithm [25], the predecessor of AURORA, also utilizes an autoencoder with RGB images as input to construct a low-dimensional behavior space, or outcome space as they characterize it. The authors, also use two metrics to decide the fitness of a solution, the fitness score and the autoencoder’s reconstruction error, and one of them is chosen for each iteration with a probability of 50%. Similarly, techniques leveraging deep reinforcement learning [26] and self-supervised approaches [27] have demonstrated promise in autonomous skill discovery, but their reliance on task-specific tuning can restrict generalization. Lastly, the Quality-Diversity Meta-Evolution (QD-Meta) framework [28], which evolves a population of QD algorithms to dynamically customize behavior spaces for user-defined meta-objectives. By using the Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES), QD-Meta uses neural networks that transform high-dimensional behavior features into compact descriptors while leveraging reinforcement learning for dynamic parameter control. This approach enables the creation of tailored, diverse, and high-quality solutions across various domains without relying on manually crafted behavior spaces. Our approach, VQ-Elites, builds on these advancements by combining MAP-Elites with a VQ-VAE to autonomously generate structured, and flexible behavior spaces, thereby overcoming the limitations of previous unsupervised methods.
IV Vector-Quantized Elites
The primary goal of this work is to develop a novel QD framework that enables the development of algorithms that are able to operate entirely unsupervised, eliminating the need for prior knowledge about the task, environment, or dynamic constraints. Our proposed VQ-Elites algorithm achieves this by leveraging the agent’s experiences in conjunction with Vector Quantized (VQ) machine learning models. This combination allows VQ-Elites to dynamically and autonomously discover meaningful behavioral descriptors that adapt to the specific characteristics of the problem at hand, ensuring flexibility and scalability across diverse tasks and environments. A detailed overview of the algorithm is provided in Algo. 2 and in Fig. 1.
Algorithm 2 VQ-Elites algorithm (number of iterations $N_{\text{evo}}$; model update interval $N_{\text{update}}$; training epochs $\mathcal{E}$; cooperation iterations $N_{\text{cooperation}}$)
1:Define: $\mathcal{M},\;\mathcal{C}$triangleright VQ-VAE model, behavioral space grid
2:Initialize: $\mathcal{D}\leftarrow\varnothing$, $\mathcal{M}_{\text{codebook}}$triangleright Initialize an empty dataset, Initialize model’s codebook randomly or using K-Means
3:procedure VQ-Elites
4:     $\mathcal{A}\longleftarrow\{\varnothing\}_{|C|}$triangleright Initialize an empty archive of size $|C|$ (number of cells in the grid)
5:     Bootstrap()
6:     for $i=1\to I$ do
7:         Evolution Phase() $\triangleright$ Run the evolution phase in every iteration
8:         if $i\mod N_{\text{update}}=0$ then $\triangleright$ Update the model every $N_{\text{update}}$ iterations
9:              UpdateModel($\mathcal{D}$) $\triangleright$ Update the model
10:         end if
11:     end for
12:end procedure
13:procedure Bootstrap
14:     $\mathcal{D}_{\text{exploration}}\leftarrow\pi_{\theta}(\alpha|s)_{K}$triangleright Explore the environment $K$ times using random policies and store the data
15:     $\mathcal{M}$, $\mathcal{D}_{\text{exploration}}$) $\triangleright$ Train the VQ-VAE model using the exploration data
16:     $\mathcal{C}\leftarrow\mathcal{M}_{\text{codebook}}$triangleright Set the learned codebook as the initial BD grid
17:     $\mathcal{A}\leftarrow\mathcal{M}_{\text{encoder}}(\mathcal{D}_{\text{exploration}})$triangleright Add random solutions to the archive
18:end procedure
19:procedure Evolution Phase
20:     $\boldsymbol{\theta}=$ selection($\mathcal{A}$) $\triangleright$ Select randomly solutions from the archive
21:     $\boldsymbol{\theta}^{\prime}=$ variation($\boldsymbol{\theta}$) $\triangleright$ Perform mutation and crossover
22:     $(fitness,d)\longleftarrow$ evaluate($\boldsymbol{\theta}^{\prime}$) $\triangleright$ Evaluate the solution, and get the fitness and the raw data
23:     $\mathcal{D}^{(i)}\longleftarrow d$triangleright Add the raw data to the training dataset
24:     $\mathbf{bd}\longleftarrow\mathcal{M}_{encoder}(d)$triangleright Extract the behavioral descriptor
25:     $c\longleftarrow$ get_cell_index($\mathbf{bd}$, $\mathcal{C}$) $\triangleright$ Get the index of the cell in the grid the bd is assigned to
26:     if $\mathcal{A}(c)=null$ or $\mathcal{A}(c).fitness<fitness$ then $\triangleright$ Check if cooperation mode is on or if the cell is occupied and compete
27:         $\mathcal{A}(c)\longleftarrow fitness,\boldsymbol{\theta}^{\prime}$triangleright Add solution to the archive
28:     end if
29:end procedure
30:procedure UpdateModel($\mathcal{D}$)
31:     $\mathcal{M}$, $\mathcal{D}$) for $\mathcal{E}$ epochs $\triangleright$ Update the VQ-VAE model
32:     $\mathcal{C}\leftarrow\mathcal{M}_{\text{codebook}}$triangleright Set the the updated codebook as the BD grid
33:     update_archive($\mathcal{A}$, $\mathcal{M}_{\text{encoder}}$) $\triangleright$ Update the BDs of the individuals in the archive, and reassign them to the closest cell
34:end procedure
A defining feature of the VQ-Elites algorithm is its unsupervised nature, which allows it to autonomously learn behavioral descriptors for the discovered solutions. Unlike traditional QD algorithms, where users must manually define BDs [3, 6] or rely on prior knowledge to generate BDs using clustering methods [18], VQ-Elites removes these limitations. Manual or predefined BD selection can restrict the range of problems the algorithm can address or render it impractical in some cases. To overcome this, VQ-Elites leverages representation learning and unsupervised methods to learn BDs directly from the data.
In this work, we employ VQ-VAEs, a generative model that combines autoencoders with vector quantization to create discrete latent representations. VQ-VAEs learn a codebook of vectors, mapping continuous latent vectors onto discrete codes. For high-dimensional data such as images, videos, or multivariate time-series, this mapping typically occurs along the channel dimension. However, when applied to one-dimensional data, VQ-VAE performs gradient-based clustering of the latent vectors, where the discrete representations correspond to centroids — similar to the K-Means algorithm [29].
This clustering property aligns seamlessly with MAP-Elites and the QD framework in general, as it provides both a mechanism for extracting BDs by encoding solution behaviors and a discrete grid (the codebook) to populate with elite solutions. Unlike classical clustering methods, this data-driven approach adapts more effectively to the dynamic nature of the problem, making VQ-Elites significantly more versatile than traditional QD methods.
Alternative unsupervised methods for BD extraction, such as AURORA [9], face significant challenges when learning from online QD data. The dynamic nature of QD archives means that the behavioral space evolves continuously, making it difficult for these methods to adapt effectively in real-time. Moreover, AURORA relies on additional hyperparameters, such as the minimum distance threshold for storing solutions in its unstructured archive. These hyperparameters can be difficult to tune and may require prior domain knowledge, reducing the method’s generality and scalability. While AURORA can be effective in certain scenarios, its reliance on unstructured archives and extra hyperparameters makes it less suitable for tasks that benefit from a structured, adaptive behavior space like the one provided by MAP-Elites.
VQ-Elites is split into two phases:
  • Bootstrap phase: in this phase, we bootstrap the archive and the unsupervised learning process via random policies;
  • Evolution phase: this is the normal operation of the VQ-Elites algorithm, where every $N_{\text{update}}$ iterations we update the unsupervised model.
IV-A Bootstrapping
Initially, the codebook is randomly initialized, something that can slow down the convergence of our algorithm. To overcome this, we first explore the environment to collect some data with random policies. Then, we use the acquired data to bootstrap our codebook by training our model a sufficient amount of epochs, e.g. $\mathcal{E}=100$. Through this process, we basically warm-start our model and create an initial grid for the evolution phase. In our experience, this step is not mandatory, but it can speed up the convergence of VQ-Elites considerably. Lastly, even though we have not investigated this direction, the user could bootstrap the model with prior information different than exploration data or use a pre-trained model instead.
IV-B Evolution Phase
The next key component of VQ-Elites is the evolution phase. During this phase the Quality-Diversity algorithm is executed, performing the usual steps. Foremost, the archive is randomly initialized with random solutions. Then, a random batch of solutions is selected, which are alternated through mutation and crossover. The new variation of the solutions are executed, and evaluated through their behavioral descriptor. In order to extract these BDs, we encode the raw data produced by our solutions with the encoder of our VQ-VAE model, acquiring the corresponding continuous latent vectors. For each BD, we find the closest grid cell, and if the cell is empty the solution is added to the archive, otherwise, we compare the fitness of the current solution with the one of the solution that is already in the archive. The solution in the archive is replaced only if the new solution is fitter. This procedure is repeated for a predefined number of iterations. Another important aspect of our framework is that the VQ-VAE model and the MAP-Elites algorithm are completely independent. Hence, the user has the freedom to incorporate different variations of the MAP-Elites algorithm as well as other selection, mutation, and crossover operators.
Updating the model: The final component of our framework is the model update, refining its latent space and codebook. The update of the model occurs every $N_{\text{update}}$ iteration and in an online fashion, meaning that the model never resets. Ultimately, the model will converge, and thus, there will be no further changes in the latent space and codebook. To train the model, we exploit the experiences of our agent, by storing the raw data produced by the solutions in the evolution phase. By utilizing these experiences, we can achieve a latent space that encapsulates the underlying dynamics of our agent as well as of the environment. Hence, the behavioral descriptors and the grid of our behavioral space consist only of feasible behaviors to be produced by the agent, something that would have been impossible to achieve by using a naive clustering algorithm.
After each update of the VQ-VAE model, we retain a new behavioral space, which foreshadows that our current grid and BDs are not aligned. Specifically, this means that the same solution before and after the update of the model is not described by the same BD. To correct this, we need to also update our archive and grid of the behavioral space. Consequently, we set the updated codebook as our new grid, and then we reevaluate the solutions we have in the archive, by encoding their raw data and assigning their BDs to the the closest cell as in the evolution phase. The update of the archive serves a double purpose; our solutions are accurately represented by the behavioral space, and both unfit and less diverse solutions are discarded due to the new refined latent space and codebook.
IV-C Behavioral Space Bounding
Having discussed the backbone of the VQ-Elites algorithm, we will introduce two key components, which ensure that QD algorithms and unsupervised learning can be coupled effectively, producing high-quality and diverse solutions while performing robustly across tasks. At the core of our approach is the manifold hypothesis, which underpins most representation learning models. This hypothesis suggests that high-dimensional datasets lie on low-dimensional latent manifolds within the high-dimensional space. As a result, data points with similar underlying information are clustered closely in the latent space, while dissimilar data points are farther apart.
While this property is advantageous in many contexts, it introduces a potential issue in our framework. Recall that our goal is to learn a behavioral space, represented by a latent space, that allows us to find solutions that are both diverse and high-performing. According to the manifold hypothesis, behavioral descriptors of diverse solutions will naturally be far apart in the latent space. However, the VQ-VAE model, by design, updates its codebook less effectively than the latent space itself. This disparity can result in an archive filled with high-performing solutions that lack diversity. Additionally, training the model in an online manner also introduces a lot of difficulties, since, until the model converges, newly found diverse solutions are actually out-of-distribution, likely leading to their latent representation not being representative.
To address this challenge, we constrain the latent (and therefore the behavioral) space to $[-1,1]$. This bounded space ensures that the latent space remains continuous and expressive, while preventing excessive dispersion that might hinder the codebook’s ability to form a meaningful discretization of the behavioral space. By maintaining this bounded structure, we enable the model to produce a diverse set of solutions while preserving high performance.
With this bounded latent space in place, we can further exploit our model to accelerate convergence. Initially, both the VQ-VAE model and its codebook are randomly initialized, which typically requires a significant number of gradient steps to achieve a well-structured discretization. However, given the fixed bounds of the latent space ($[-1,1]$), we can initialize the codebook more effectively. Specifically, we pre-train the codebook by fitting the K-Means algorithm on a sufficient number of samples drawn from the uniform distribution, $\mathcal{U}_{[-1,1]}$.
This initialization provides the codebook with centroids that are already representative of the bounded latent space. As a result, the model requires fewer gradient updates to achieve an optimal discretization of the behavioral space. This accelerates both the learning process and the creation of a well-structured archive, ensuring that the algorithm efficiently identifies diverse, high-quality solutions.
IV-D Cooperation over Competition
During the evolution phase, new solutions are introduced to the archive by competing with existing ones. However, in the early iterations, the archive often contains less diverse solutions because the representation learning model has not yet fully established a well-defined behavioral space. As a result, early individuals in the archive may achieve high fitness scores but lack diversity.
To address this issue, we introduce a cooperation aspect to our algorithm during the initial iterations. This cooperation neglects the competition of individuals, and instead, the grid is filled with the most recent solution assigned to each cell, regardless of fitness. In a way, the individuals are cooperating with offsprings of the population, with the goal of promoting the diversity of the archive. This approach allows the model to focus on constructing a highly diverse grid or refining the behavioral space before strict competition begins. Additionally, this strategy enhances the algorithm’s exploration capabilities by ensuring frequent updates to the solution set, promoting diversity during the critical early stages of the evolutionary process. We should emphasize that the VQ-Elites is robust and did not require the cooperation component for our experiments, yet we found it critical for other unsupervised QD algorithms we evaluated.
V Experiments
For the experimental part of our proposed approach, we want to include a representative set of experiments that answer the following questions:
  1. 1.
    How does VQ-Elites compare to the classic MAP-Elites when we have full knowledge of the environment?
  2. 2.
    How does VQ-Elites performs when we do not possess prior knowledge about the environment or the constraints of the agent?
  3. 3.
    Is VQ-Elites robust for a range of hyperparameters or requires extensive fine-tuning?
  4. 4.
    How vital are the behavior space bounding and cooperation components in the convergence of unsupervised QD algorithms?
For our experiments, we selected two state-covering tasks using a mobile robot and two pose-reaching tasks with a 7 degrees of freedom robotic arm. These tasks were chosen to showcase the versatility and effectiveness of VQ-Elites in solving robotics problems without requiring prior knowledge of the environment or the system. Moreover, these experiments demonstrate the broader potential of VQ-Elites beyond these specific scenarios.
A key advantage of VQ-Elites is its minimal reliance on hyperparameter tuning. Throughout all experiments, the crossover and mutation rates were kept consistent, and the model was updated every five iterations of the evolutionary process. This consistency highlights the robustness and ease of deployment of VQ-Elites across diverse tasks.
V-A Mobile Robot Experiments
For the space-covering experiments, we employ a differential drive mobile robot. The robot is deployed in two environments, one without obstacles and one with a L-shaped environment as depicted in Fig. 2. During the evolution phase, the algorithm evolves a population that contains the weights of a NN, which acts as the policy. This policy takes as input the state of the robot $\mathbf{s}=\{x,y,\cos(\theta),\sin(\theta)\}$, and produces a $\mathbf{v}=\{v_{L},v_{R}\}$ that includes the velocity that should be applied on each wheel. The goal of this task is for the robot to cover the whole environment while reaching each final position as fast as possible. Hence the fitness function is:
$r=-\sum_{i}^{N}||\boldsymbol{p}_{i}-\boldsymbol{p}_{N}||_{2},$ (8)
where $\boldsymbol{p}_{i}$ is the position of the robot in the $i$-th step of its trajectory. For every candidate solution, the raw BD is an image of the last position of the robot in the environment with dimensions $64\times 64$, which we feed into the encoder network of VQ-VAE to extract our low-dimensional BD. Additionally, these images are stored for the learning phase of our pipeline. We should note, that the reward of the environment is only responsible for the fitness of the solution. The space-covering aspect of the task is fulfilled in the learning phase, and thus, the model should construct an optimal desensitization of our behavioral space to achieve it.
During the learning phase, we noticed that the model became biased due to the constructed dataset being unbalanced. This phenomenon arises as a combination of two things, the ratio of important information in the images is low and at the beginning of our algorithm, the solutions are not diverse enough. To overcome that, we filter our dataset by removing instances that are very similar to each other, by checking the amount of overlap the robot has on the pixel level.
Lastly, the initial position of the robot is fixed, and the algorithm is executed for 3000 iterations, while each solution performs 400 steps in the environment.
Figure 2: Visualization of the obstacle-free and maze environments for the mobile robot experiments.
V-B Robotic Arm Experiments
Regarding the pose-reaching tasks, the general notion of the pipeline follows the mobile robot experiment, with some variations. The goal of the task is for the end-effector of the robot to reach a predefined location, with as diverse joint configurations as possible. Here, the input of our policy as well as the data we use to extract the BD and train the VQ-VAE model are the joints of the robot (raw BD). The output of the policy is the joint velocities.
In order to demonstrate the flexibility of our algorithm, we first showcase this experiment with the default limits of the joint space, and after, we introduce some artificial constraints to the robot, which simulate a problematic agent or a confined operation space. The goal of this variation is to exhibit the high potential of unsupervised learning in combination with QD algorithms in solving problems in an agnostic manner, where the constraints of the system cannot be identified beforehand. For both of these tasks, we use 3000 iterations and each policy performs 300 environment steps.
V-C Evaluation Metrics
To evaluate the performance of our algorithm, we employ most of the metrics used in AURORA [9]. Specifically, we evaluate our approach in terms of coverage, grid mean fitness, and QD score.
V-C1 Coverage
This metric measures the percentage of the hand-coded behavioral space that is covered. The BD space is divided into a grid of cells, and the coverage is calculated as the proportion of cells containing at least one individual. In the mobile robot experiments, since we operate in a 2D space, we simply discretize the robot’s operating space in cell of dimension $N\times N$. On the other hand, this is not possible in the pose-reaching task, hence, we gather 1M poses using inverse kinematics (IK) and perform a K-Means clustering on these data. The resulting centroids correspond to our hard-coded BD.
V-C2 Grid Mean Fitness
This metric assesses the overall quality of a container by calculating the mean fitness of individuals in a $N\times N$ grid of the hand-coded BD space. For each cell, only the individual with the highest fitness is retained, and their fitnesses are averaged to compute the grid mean fitness. This approach ensures that the metric is independent of the solution distribution within the BD space, enabling fair comparisons across algorithms using learned BD spaces.
V-C3 Quality-Diversity Score
The QD score is used to assess the performance of an algorithm in terms of both the diversity and the quality of solutions in the BD space. It is calculated as the sum of the fitness values of all the individuals across the grid. By combining diversity (coverage of the grid) and quality (fitness of solutions), the QD score provides a single, unified measure to evaluate the effectiveness of a given algorithm. Higher QD scores indicate better exploration and exploitation of the BD space.
TABLE I: List of all the hyperparameters used for each type of environment
      Experiment Hyperparameters
      Experiment       Mobile robot       Robotic arm
      Iterations       $3\times 10^{3}$       $3\times 10^{3}$
      Population size       $128$       $128$
      Mutation rate       $10\%$       $30\%$
      Cross-over rate       $10\%$       $20\%$
      $|\mathcal{C}|$       $2000$       $400$
      BD/Latent size       $2$       $5$
      $\mathcal{E}$       $10$       $10$
      $lr$       $7\times 10^{-4}$       $7\times 10^{-4}$
      Batch size       $64$       $64$
      $N_{\text{update}}$       $5$       $5$
      $N_{\text{cooperation}}$       $1\times 10^{3}$       $1\times 10^{3}$
VI Results
In this section, we showcase the results of our algorithm in the experimental evaluation, starting with the mobile robot experiments, then continuing to robotic arm experiments and finally presenting an ablation study we performed on our algorithm.
VI-A Mobile Robot Results
For the mobile robot experiments, we compare VQ-Elites algorithm with the classic MAP-Elites algorithm, the AURORA approach as well as the AURORA after integrating our two proposed components, behavioral space bounding and cooperation, which from now on we will refer as AURORA. We should mention that a variation of this experiment is also present in the AURORA paper [9], hence it is ideal for comparison purposes, however, we are unable to exactly replicate the authors’ results [^1] We used the exact hyperparameters presented in the original paper and communicated with the authors., as demonstrated in the manuscript.
In the case where the robot operates in an obstacle-free environment, we can observe that all the algorithms, except the vanilla AURORA, perform almost perfectly in terms of coverage, assigning solutions to nearly every grid cell (Fig. 3). In terms of the QD score, we see that the classic MAP-Elites algorithm produces more high-quality and diverse solutions compared to the other approaches. Both of these observations are expected, since we can easily define a behavioral grid manually, and still every cell will represent a feasible behavior. Nonetheless, it is evident that algorithms that assume no prior knowledge about the environment such as ours and other unsupervised QD algorithms infused with our cooperation and behavioral space bounding components, can achieve the performance of MAP-Elites without full knowledge of the task (answers to Question 1).
Figure 3: Results for the mobile robot operating in the obstacle-free environment of four different algorithms, VQ-Elites, AURORA, AURORA, and MAP-Elites with hard-coded behavioral descriptors. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
Transitioning to the L-shaped environment, our approach and AURORA outperform the others by a large margin in terms of coverage, while VQ-Elites exhibits the best QD score of all (Fig. 4). This is one type of task that makes the existence of unsupervised QD algorithms essential, since, if we do not possess full knowledge about the task at hand, and particularly in this scenario, we do not know the environment, we cannot construct a grid such that every cell corresponds to a possible behavior (answers to Question 2). On the other hand, a data-driven unsupervised approach is more than capable of constructing such a grid.
Figure 4: Results for the mobile robot operating in the L-shaped environment of four different algorithms, VQ-Elites, AURORA, AURORA, and MAP-Elites with hard-coded behavioral descriptors. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
Lastly, since the experiment is two-dimensional, we also provide plots, which are hard or infeasible to do or comprehend in high dimensions. These plots depict the grid mean fitness, final solutions/archive of our algorithm, and the behavior descriptors and behavioral grid for each scenario in Fig. 5. In the left plot, we illustrate the grid mean fitness by discretizing the robot’s operating space into a $8\times 8$ grid, and the color of each cell represents the normalized reward/fitness of the best solution assigned to each cell. The middle plot showcases the actual solutions VQ-Elites produced, while the right plot depicts the BDs of the solutions that are shown in the middle plot along with all the centers of our behavioral space grid learned by the VQ-VAE model. Overall, we observe that VQ-Elites is able to effectively explore the behavior space, produce meaningful latent representations, and automatically perform clustering of the elite behaviors.
Figure 5: Plots of the mean fitness in an 8x8 grid, final archive of solutions, and the behavior descriptors and learned behavioral grid of: (a) obstacle-free environment. (b) L-shaped environment.
VI-B Robotic Arm Results
For the robotic arm experiments, we require the robot to achieve diverse poses, filling the whole operation space. In Fig. 8, we provide a visualization of some diverse poses our algorithm achieves and the final result, where the operation space is filled. For this task, we investigate two different aspects, firstly, how the structured grid archive VQ-Elites uses compares to the unstructured archive of AURORA. Secondly, if the robotic arm system has some constraints, e.g. due to motor malfunction, and thus some joints are not fully operational, can our algorithm still find optimal solutions, and how does it compare to the classic MAP-Elites algorithm?
For the first case, we find that the coverage of VQ-Elites and MAP-Elites is identical while the AURORA performs slightly worse, yet, we can see a significant difference in terms of QD score, without a trend of improvement (Fig. 6). This fact indicates that the way an unstructured archive is updated is highly likely to discard high-quality and diverse individuals in order to satisfy the minimum distance between individuals rule (answers to Question 1).
The second case resembles the mobile robot in the L-shaped environment. However, in this task, the constraint is on the robot itself. Despite that, our algorithm has a completely agnostic nature due to the unsupervised learning model, and hence, it can outperform the MAP-Elites algorithm (Fig. 7). We should note that this is highly predictable because in order for the MAP-Elites to achieve a competitive performance it requires knowledge of the constraints to construct a proper behavioral descriptor grid. Lastly, AURORA performs similarly to the MAP-Elites algorithm despite its unsupervised nature, while the QD score is the worst of the three due to the way the archive is constructed, as noted previously (answers to Question 2).
Figure 6: Results for the robotic arm pose reaching task for VQ-Elites, MAP-Elites, and AURORA. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
Figure 7: Results for the robotic arm pose reaching task with system constraints (i.e. limitation in some joints) for VQ-Elites, MAP-Elites with hard-coded BDs, and AURORA. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
Figure 8: Top: Illustration of a subset of diverse poses (individuals) achieved from the VQ-Elites algorithm. Botton: The final VQ-Elites archive that fills the behavior space with diverse poses.
VI-C Ablation Study
In order to identify whether VQ-Elites is robust to hyper-parameters we perform some ablation studies. In particular, we want to investigate how the combination of the model’s update frequency, the training epochs, and the use of the cooperation component, affect the performance of our algorithm. For these experiments, we used some extreme values for the mentioned hyperparameters to investigate our algorithm’s robustness. We perform the experiments in the L-shaped mobile robot environment. The results showcase that when we update the model after every iteration of the evolution phase and train the model for 1 epoch, the performance is suboptimal despite the use of the cooperation component (Fig. 9). This is anticipated since the behavioral space grid changes continuously, but also the model fails to converge fast enough due to the low number of gradient updates. Conversely, when we update the model every 100 iterations for 100 epochs, the performance is similar to our original hyperparameters (answers to Question 3). Although one could argue that this is more selection of hyperparameters compared to the ones we use in our main experiments, we should emphasize that updating the model after every 100 iterations raises memory and wall-time issues. For every iteration, one should store all the raw data produced by the populations, which can easily lead to reaching that maximum memory capacity, especially when working with high-dimensional data like images, while slowing down significantly the model’s update.
Furthermore, we investigate the importance of the cooperation component on our algorithms as well as on the AURORA algorithm. As stated, VQ-Elites is quite robust that the cooperation does not have any impact on its performance, since the VQ-VAE model, eventually, coverages to an optimal grid structure (answers to Question 4). On the other hand, the AURORA algorithm performs poorly without this component (Fig. 10).
Figure 9: Results of the ablation study regarding the model’s update frequency, the number of training epochs, and the use of cooperation component on the mobile robot maze experiment. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
Figure 10: Illustration of the effect of the cooperation component on VQ-Elites and AURORA.
VII Conclusion and Discussion
In this work, we presented the Vector Quantized-Elites, VQ-Elites, algorithm, an unsupervised and problem-agnostic variation of the MAP-Elites algorithm. Our algorithm requires no prior knowledge of the task at hand to provide high-quality and diverse solutions. Additionally, we proposed two key components, the behavioral space bounding and cooperation components, which can be integrated into other unsupervised QD algorithms to mitigate the difficulties between the evolution and the update of the unsupervised model. Our work is deployed on two robotic systems, a mobile robot, and a robotic arm, while its performance is evaluated on numerous experiments. In all scenarios, our algorithm performs as well as or better than the AURORA algorithm (even when coupled with our two novel components, AURORA), making VQ-Elites algorithm a strong candidate to tackle many optimization problems.
Nevertheless, our work does not come without its limitations. It is important to note that our approach can become very costly in terms of computation time or exhibit degraded performance. The main cause for these problems originates from the selection of the raw data and the unsupervised model’s architecture. For example, utilizing very inefficient architectures for the VQ-VAE or using memory-intensive raw data (e.g. videos) to extract the BDs as well as train the model can only hinder our approach’s efficiency. Additionally, given the raw data, the selection of the model’s architecture should be tuned since a non-optimal model will not be able to construct a meaningful latent space, and consequently, a high-quality behavioral grid, resulting in poor solutions for the task at hand.
In the future, we want to incorporate generative modeling techniques such as diffusion, or flow-matching, in order to model the distribution of solutions and generate more high-quality and diverse solutions. An ideal outcome would be to use VQ-Elites for finding the minimum amount of solutions, and then via the generative modeling techniques, we would able to fill the archive successfully without being limited by the run times of the evolutionary algorithm and its sequential nature. Lastly, another interesting direction would be an epistemic uncertainty-guided evolution frame. The intuition here lies in the fact that epistemic uncertainty decreases when the number of data we train our model increases, hence, we can use the epistemic uncertainty of the model as a way to guide the QD evolution into more uncertain behaviors, with the goal of reducing the epistemic uncertainty of the VQ-VAE model, and consequently creating the most optimal behavioral space possible.
Acknowledgments
This work was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I. Research Projects to support Post-Doctoral Researchers” (Project Acronym: NOSALRO, Project Number: 7541).